Find, showing all working, a formula for the n-th term tn of the sequence (tn) defined by t1 = 5; tn = −7tn−1/3, n ≥ 2.


1
Expert's answer
2021-10-12T11:17:03-0400
"r=\dfrac{t_n}{t_{n-1}}=\dfrac{-\dfrac{7t_{n-1}}{3}}{t_{n-1}}=-\dfrac{7}{3}, n\geq2"

We have a geometric progression


"a=t_1=5, r=-\dfrac{7}{3}"

"t_n=5(-\dfrac{7}{3})^{n-1}, n\geq2"